the primeter of a triangle is 63cm. The lengths of the three sides are represented by three consecutive odd integers. Find the length of the longest side
Answer by checkley71(8403)

X+X+2+X+4=63
3X+6=63
3X=63-6
3X=57
X=57/3
X=19 IS THE SHORTEST SIDE
19+4=23 IS THE LONGEST SIDE
PROOF
19+21+23=63
63=63